OpenStudy (anonymous):

subtract. simplify, if possible(be careful) x + 2 ---- ----- x^2+5x+6 x^2+{3x+2

jhonyy9 (jhonyy9):

x(x*2 -3x +2) + 2x*2 +10x +12 =x*3 -3x*2 +2x +2x*2 +10x +12 = =x*3 -x*2 +12x +12